OpenStudy (anonymous):

Audrey is trying to find the equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1). Which of the following equations will she use?

OpenStudy (anonymous):

y − (−6) = 2 over 3(x − (−1)) y − (−1) = 2 over 3(x − (−6)) y − (−6) = 3 over 2(x − (−1)) y − (−1) = 3 over 2(x − (−6))

OpenStudy (anonymous):

its either a or c

OpenStudy (anonymous):

It is a

OpenStudy (anonymous):

for sure

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

will u medal me plz?